State the domain and range of f(x)=2^x. Here is what I did.....The domain of f(x)=2^x would be the x values. This would include all values that you can input as x in order to make this problem work. The domain of a function is usually all real numbers. The range of f(x)=2^x would be the y values. This would include all values that would be the output for the y value. An example of this would be if you used 2 as x then the function would read f(x)=2^2. The y would equal 4 which would be included in the range of this function. To find the domain and range of the inverse you would follow the proper steps to get the inverse of the function which would be x=2^y. The domain would be the x values and the range would be the y values. If you put 4 as x which would be your input for the domain you would get 2 for the y which would be the range.
I understand that the domain is all real numbers, but I still have to be able to state the entire range of the function. Thanks!

Answer by stanbon(75887)

State the domain and range of f(x)=2^x. Here is what I did.....The domain of f(x)=2^x would be the x values. This would include all values that you can input as x in order to make this problem work. The domain of a function is usually all real numbers. The range of f(x)=2^x would be the y values. This would include all values that would be the output for the y value. An example of this would be if you used 2 as x then the function would read f(x)=2^2. The y would equal 4 which would be included in the range of this function. To find the domain and range of the inverse you would follow the proper steps to get the inverse of the function which would be x=2^y. The domain would be the x values and the range would be the y values. If you put 4 as x which would be your input for the domain you would get 2^4 = 16 for the y which would be the range.
I understand that the domain is all real numbers, but I still have to be able to state the entire range of the function. Thanks!
---------
y = 2^x
----
Domain: All Real Numbers
Range: 2^x can never be 0 or negative
So, Range is All Real Numbers greater than zero.
-------------------
Cheers,
Stan H.
------------